eral Physics 1 & 2 (Physics 181/182) Chapter 01, Problem 23 Two workers are trying to move a heavy crate. One pushes on the crate with a force pushes with a force B, which has a magnitude of 442 N and is directed due north. whet are (a) the magsitude and (b) direction of the resultant force 2 applied to the erate? Suppose that the second worer spoies aforce .Bmered o , what then ro (c)the ina-de and Cndeden of the r-rt hre ajapeled the crate? In both cases A. which has a magnitude of 216 newtons (N) and is directed due west. The other express the direction as a positive angle relative to due west (a) Number (b) Number Donh of west (e) Number Id) Number south of Question Attempts: Unlimited VE POR LATER 3 5 6 8 9 commandoption
Answer #1

A)

|A + B| = sqrt(A^2 + B^2) = sqrt(216^2 + 442^2) = 491.96 N

Theta = arctan (442/216) = 63.96 degree (North of west)

B)

|A - B| = sqrt(216^2 + 442^2) = 491.96 N

Theta = arctan (-442/216) = 63.96 degree (South of west)

Comment in case any doubt please rate my answer ....